# bibliography.bib

### v1

@@ 5 Canonical extensions as Raney extensions
If $X$ is strictly sober, $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{SO}$-compact, by Lemma 5.4. It is also $\mathcal{CP}$-dense, by Lemma 3.10, and as completely prime filters are Scott-open it is also $\mathcal{SO}$-dense. Conversely, if $X$ is a space such that $(\Omega(X),\mathcal{U}(X))$ is a canonical extension, in particular this Raney extension is $\mathcal{SO}$-compact, hence $X$ is strictly sober by Lemma 5.4. ∎
It is known that if we do not assume choice principles it is not the case that sobriety implies strict sobriety. Let us look at a counterexample for this.
@@
We assume the negation of the Ultrafilter Lemma and deduce that there exists a sober space which is not strictly sober. Let $X$ be a set and let $\mathcal{P}(X)$ be its powerset, let $F\subseteq\mathcal{P}(X)$ be a filter such that it is not contained in any ultrafilter. Now, consider the Stone dual $X^{S}$ of $\mathcal{P}(X)$, let $\varphi$ be its topologizing map. Note that this is an isomorphism of Boolean algebras, as $\mathcal{P}(X)$ is atomic. All elements of the form $\varphi(Y)$ for $Y\subseteq X$ are clopens of the space $X^{S}$, hence compact. We then have that the filter of opens ${\uparrow}\varphi[F]$ is then Scott-open. By assumption, $\bigcap\varphi[F]=\emptyset$, and so if ${\uparrow}\varphi[F]$ is a neighborhood filter of some compact open this must be $\emptyset$, but this is not the case as the neighborhood filter of $\emptyset$ contains $\emptyset$, and $\emptyset\notin\varphi[F]$ by injectivity of $\varphi$. Thus, the space $X^{S}$ is sober, as it is a Stone space, but it is not strictly sober.
@@
The Prime Ideal Theorem is equivalent to the statement that every Scott-open filter is an intersection of completely prime filters.
@@
    We have $\mathsf{Filt}_{\mathcal{SO}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$ for every frame $L$.
@@
Finally, we give another proof, based on collections of filters, of the result in [bezhanishvili22] that the canonical extension of a Boolean algebra $B$ is the Booleanization of $\mathcal{U}(\mathsf{Idl}(B))$.
@@
We have that $\neg{\uparrow}k=\{a\in L:a\vee k=1\}$, and this equals ${\uparrow}\neg k$. By definition of complement, also ${\uparrow}k\cap{\uparrow}\neg k=\{1\}$ and ${\uparrow}k\vee{\uparrow}\neg k=L$. ∎
@@
Let $L$ be a compact, zero-dimensional frame, and let $F$ be a Scott-open filter. Let $f\in F$. We have $F=\bigvee_{i}k_{i}$, where $\{k_{i}:i\in I\}$ is the family of complemented elements below it. Observe that this is directed, and so there must be $j\in I$ with $k_{j}\in F$. For the converse, if $K$ is any family of complemented elements, suppose that there is a directed family $D\subseteq L$ such that $k\leq\bigvee D$ for some $k\in D$. As $L$ is compact, $k$ is compact, too, and so $k\leq d$ for some $d\in D$. Then, $\bigvee_{k\in K}{\uparrow}k$ is Scott-open. ∎
@@
For a compact, zero-dimensional frame $L$, we have $\mathsf{Filt}_{\mathcal{R}}(L)=\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$.
@@
Because $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$, and this is the smallest sublocale containing $\{1\}$, for the inclusion $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$ it suffices to show that $\{1\}$ is Scott-open, but this follows immediately from compactness of $L$. For the other direction, it suffices to show that for a Scott-open filter $F$ we have $\neg\neg F\subseteq F$. Let $F$ be a Scott-open filter. By Lemma 5.10, this is $\bigvee_{i}{\uparrow}k_{i}$ for some collection $k_{i}\in L$ of complemented elements, which we can assume to be closed under finite meets without loss of generality. We have $\neg\neg\bigvee_{i}{\uparrow}k_{i}=\neg\bigcap_{i}\neg{\uparrow}k_{i}=\neg\bigcap_{i}{\uparrow}\neg k_{i}$, where we have used Lemma 5.9 for the last equality. Note also that $\neg\bigcap_{i}{\uparrow}\neg k_{i}=\neg{\uparrow}\bigvee_{i}\neg k_{i}$. Now, if $x\in\neg\neg F$ this means that $\bigvee_{i}\neg k_{i}\vee x=1$, and by compactness this means that $\neg k_{j}\vee x=1$ for some $j\in I$. Therefore $k_{j}\leq x$, and so $x\in F$. ∎
@@
The first part of the claim follows from Lemma 5.11. The second part of the claim follows from Proposition 5.8, and the fact that $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$. This is also the Booleanization of $\mathcal{U}(L)$, because for each $x\in L$ we have $\neg{\uparrow}x=\{y\in L:x\vee y=1\}$, and so the regular elements of $\mathcal{U}(L)$ are precisely the intersections of upsets of this form, that this, the regular filters of $L$. ∎

### v2

@@ 7.1 Sobriety and strict sobriety
If $X$ is strictly sober, $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{SO}$-compact, by Lemma 7.1. It is also $\mathcal{CP}$-dense, by Corollary 4.10, and as completely prime filters are Scott-open it is also $\mathcal{SO}$-dense. Conversely, if $X$ is a space such that $(\Omega(X),\mathcal{U}(X))$ is a canonical extension, in particular this Raney extension is $\mathcal{SO}$-compact, hence $X$ is strictly sober by Lemma 7.1. ∎
It is known that if we do not assume choice principles it is not the case that sobriety implies strict sobriety. Let us look at a concrete counterexample for this.
@@
We assume the negation of the Ultrafilter Lemma and deduce that there exists a sober space which is not strictly sober. Let $X$ be a set and let $\mathcal{P}(X)$ be its powerset, let $F\subseteq\mathcal{P}(X)$ be a filter such that it is not contained in any ultrafilter. Now, consider the Stone dual $X^{S}$ of $\mathcal{P}(X)$, let $\varphi$ be its topologizing map. Note that this is an isomorphism of Boolean algebras, as $\mathcal{P}(X)$ is atomic. All elements of the form $\varphi(Y)$ for $Y\subseteq X$ are clopens of the space $X^{S}$, hence compact. We then have that the filter of opens ${\uparrow}\varphi[F]$ is Scott-open. By assumption, $\bigcap\varphi[F]=\emptyset$, and so if ${\uparrow}\varphi[F]$ is a neighborhood filter of some compact open. This must be $\emptyset$, but this is not the case as the neighborhood filter of $\emptyset$ contains $\emptyset$, and $\emptyset\notin\varphi[F]$ by injectivity of $\varphi$. Thus, the space $X^{S}$ is sober, as it is a Stone space, but it is not strictly sober.
We want to rephrase the SPET property in terms of Raney extensions.
@@
The Prime Ideal Theorem is equivalent to the statement that $\mathsf{Filt}_{\mathcal{SO}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$ for every frame $L$.
@@
    $\mathsf{Filt}_{\mathcal{SO}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$ for every frame $L$.
@@
Finally, we give another proof, based on Raney extensions, of the result in [bezhanishvili22] that the canonical extension of a Boolean algebra $B$ is the Booleanization of $\mathcal{U}(\mathsf{Idl}(B))$.
@@
Since $\neg{\uparrow}k=\{a\in L:a\vee k=1\}$, $\neg{\uparrow}k={\uparrow}\neg k$. By definition of complement, also ${\uparrow}k\cap{\uparrow}\neg k=\{1\}$ and ${\uparrow}k\vee{\uparrow}\neg k=L$. ∎
@@
Let $L$ be a compact, zero-dimensional frame, and let $F$ be a Scott-open filter. Let $f\in F$, and let $\{k_{i}:i\in I\}$ be the family of complemented elements below it, Then, $F=\bigvee_{i}k_{i}$. Observe that this is directed, and so there must be $j\in I$ with $k_{j}\in F$. For the converse, if $K$ is any family of complemented elements, suppose that there is a directed family $D\subseteq L$ such that $k\leq\bigvee D$ for some $k\in D$. As $L$ is compact, $k$ is compact, too, and so $k\leq d$ for some $d\in D$. Then, $\bigvee_{k\in K}{\uparrow}k$ is Scott-open. ∎
@@
For a compact, zero-dimensional frame $L$, $\mathsf{Filt}_{\mathcal{R}}(L)=\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$.
@@
Because $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$, and this is the smallest sublocale containing $\{1\}$, for the inclusion $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$ it suffices to show that $\{1\}$ is Scott-open, but this follows immediately from compactness of $L$. For the other direction, it suffices to show that, for a Scott-open filter $F$, $\neg\neg F\subseteq F$. Let $F$ be a Scott-open filter. By Lemma 7.7, this is $\bigvee_{i}{\uparrow}k_{i}$ for some collection $k_{i}\in L$ of complemented elements, which we can assume to be closed under finite meets without loss of generality. The equalities $\neg\neg\bigvee_{i}{\uparrow}k_{i}=\neg\bigcap_{i}\neg{\uparrow}k_{i}=\neg\bigcap_{i}{\uparrow}\neg k_{i}$ hold, where Lemma 7.6 is used for the last equality. Note also that $\neg\bigcap_{i}{\uparrow}\neg k_{i}=\neg{\uparrow}\bigvee_{i}\neg k_{i}$. Now, if $x\in\neg\neg F$ this means that $\bigvee_{i}\neg k_{i}\vee x=1$, and by compactness this means that $\neg k_{j}\vee x=1$ for some $j\in I$. Therefore $k_{j}\leq x$, and so $x\in F$. ∎
@@
The first part of the claim follows from Lemma 7.8. The second part of the claim follows from Proposition 7.5, and the fact that $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$. This is also the Booleanization of $\mathcal{U}(L)$, because, for each $x\in L$, $\neg{\uparrow}x=\{y\in L:x\vee y=1\}$, and so the regular elements of $\mathcal{U}(L)$ are precisely the intersections of upsets of this form, that this, the regular filters of $L$. ∎
7.2 Exactness and TD duality
Recall that the category $\mathbf{Frm}_{\mathcal{E}}$ is embedded in $\mathbf{Raney}$ as the full subcategory of $T_{D}$ objects. In this section, we study $\mathbf{Frm}_{\mathcal{E}}$ as a pointfree category of $T_{D}$ spaces.
Lemma 7.10.
Exact morphisms are D-morphisms.
Proof.
Suppose that $f:L\to M$ is an exact frame map, and let $p\in M$ be a covered prime. By Proposition 4.12, $L{\setminus}{\downarrow}p$ is exact. By exactness of $f$, so is $f^{-1}(L{\setminus}{\downarrow}p)$. By adjointness, $f^{-1}(L{\setminus}{\downarrow}p)=L{\setminus}{\downarrow}f_{*}(p)$. By Proposition 4.12 again, $f_{*}(p)$ is covered. ∎
Lemma 7.11.
Any $T_{D}$ frame map $f:L\to M$ such that $M$ is $T_{D}$-spatial is exact.
Proof.
Suppose that $L$ and $M$ are frames and $M$ is $T_{D}$-spatial, and that there is a frame map $f:L\to M$ such that $f_{*}(p)$ is a covered prime whenever $p\in M$ is covered. Now, suppose that $\bigwedge_{i}x_{i}\in L$ is an exact meet. We show $\bigwedge_{i}f(x_{i})\leq f(\bigwedge_{i}x_{i})$. Suppose that $p\in M$ is a covered prime with $f(\bigwedge_{i}x_{i})\leq p$. Then $\bigwedge_{i}x_{i}\leq f_{*}(p)$, that is, $\bigwedge_{i}x_{i}\vee f_{*}(p)=f_{*}(p)$. By exactness, $\bigwedge_{i}(x_{i}\vee f_{*}(p))=f_{*}(p)$, and by coveredness there is $i\in I$ with $x_{i}\vee f_{*}(p)=f_{*}(p)$. Then, $f(x_{i})\leq p$, which implies $\bigwedge_{i}f(x_{i})\leq p$, and by $T_{D}$-spatiality this implies $\bigwedge_{i}f(x_{i})\leq f(\bigwedge_{i}x_{i})$ as desired. Let us now show that $\bigwedge_{i}f(x_{i})$ is exact. Let $y\in M$. Suppose that $\bigwedge_{i}f(x_{i})\vee y\leq p$ for $p\in M$ a covered prime. As shown above, this means $f(\bigwedge_{i}x_{i})\vee y\leq p$. Similarly as above, we obtain $f(x_{i})\vee y\leq p$ for some $i\in I$, and so $\bigwedge_{i}(f(x_{i})\vee y)\leq p$. By $T_{D}$-spatiality, $\bigwedge_{i}(f(x_{i})\vee y)\leq\bigwedge_{i}f(x_{i})\vee y$, as desired. ∎
Theorem 7.12.
There is an adjunction $\Omega:\mathbf{Top}_{D}\leftrightarrows\mathbf{Frm}_{\mathcal{E}}^{op}:\mathsf{pt}_{D}$.
Proof.
It suffices to show that the functor $\Omega$ maps continuous maps between $T_{D}$ spaces to exact frame maps, and that the $T_{D}$ spatialization map of a frame is exact. By Lemma 7.11, it is known that the spatialization map is a $T_{D}$ morphism, so by Lemma 7.11 it is also exact. By the same Lemma, a map $f:X\to Y$ between $T_{D}$ spaces determines an exact frame map $\Omega(f):\Omega(Y)\to\Omega(X)$. ∎
This means that $T_{D}$ duality remains intact if we replace $\mathbf{Frm}_{D}$ with the subcategory $\mathbf{Frm}_{\mathcal{E}}$. The advantage of working in this category is that the definition of the morphisms does not mention points, and that all morphisms $f:L\to M$ lift to morphisms $\mathsf{S}_{\mathfrak{c}}(f):\mathsf{S}_{\mathfrak{c}}(L)\to\mathsf{S}_{\mathfrak{c}}(M)$, by the isomorphism $\mathsf{S}_{\mathfrak{c}}(L)\cong\mathsf{Filt}_{\mathcal{E}}(L)$. The situation is illustrated below, where the functor $\mathsf{S}_{\mathfrak{c}}$ is the one mapping a frame $L$ to the Raney extension $(L,\mathsf{S}_{\mathfrak{c}}(L)^{op})$.
${\mathbf{Frm}_{\mathcal{E}}}$${\mathbf{Top}_{D}}$${\mathbf{Raney_{D}},}$$\scriptstyle{\mathsf{pt}_{D}}$$\scriptstyle{\mathsf{S}_{\mathfrak{c}}}$$\scriptstyle{\mathsf{pt}_{R}}$
We now look at the notion of sublocale in $\mathbf{Frm}_{\mathcal{E}}$.
Lemma 7.13.
For a frame $L$, if a meet $\bigwedge_{i}x_{i}\in L$ is exact, then so is $\bigwedge_{i}(x_{i}\vee y)$ for all $i\in I$.
Proof.
Observe that, if $\bigwedge_{i}x_{i}$ is exact, for all $z\in L$, $\bigwedge_{i}(x_{i}\vee y\vee z)\leq(\bigwedge_{i}x_{i})\vee y\vee z\leq(\bigwedge_{i}(x_{i}\vee y))\vee z$. ∎
Proposition 7.14.
A surjective frame map $f:L\to M$ such that it preserves exact meets is exact.
Proof.
Suppose that $\bigwedge_{i}x_{i}$ is exact and that $f:L\to M$ is a frame surjection which preserves exact meets. For $u\in L$, $\bigwedge_{i}(f(x_{i})\vee f(u))=\bigwedge_{i}f(x_{i}\vee u)=f(\bigwedge_{i}x_{i}\vee u)=\bigwedge_{i}f(x_{i})\vee f(u)$. We have used Lemma 7.13 for the first equality. Since all elements of $M$ are $f(v)$ for some $v\in L$, the meet $\bigwedge_{i}f(x_{i})$ is exact. ∎
We say that a sublocale is exact if the corresponding surjection is exact. Let us call $\mathsf{S}_{\mathcal{E}}(L)$ the ordered collection of exact sublocales of a frame.
Proposition 7.15.
A sublocale $S$ is exact if and only if for every exact meet $\bigwedge_{i}x_{i}$ and, for all $x\in L$, $\mathfrak{c}(x_{i})\cap S\subseteq\mathfrak{c}(x)$ for all $i\in I$ implies that $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S\subseteq\mathfrak{c}(x)$.
Proof.
@@ 7.2 Exactness and TD duality
Remark 7.2.
We note that the result above can be generalized: a sublocale $S$ is such that $\sigma_{S}$ preserves a certain class of meets if and only if for all meets $\bigwedge_{i}x_{i}$ in that class, for all $x\in L$, that $\mathfrak{c}(x_{i})\cap S\subseteq\mathfrak{c}(x)$ for all $i\in I$ implies that $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S\subseteq\mathfrak{c}(x)$.
Proposition 7.16.
The collection $\mathsf{S}_{\mathcal{E}}(L)$ is closed under all joins, and it contains
  • All closed sublocales;
  • All open sublocales;
  • The two-element sublocales $\mathfrak{b}(p)$ for covered $p$.
Proof.
By Proposition 7.15, if $S_{j}$ is a collection of exact sublocales, and $\bigwedge_{i}x_{i}$ an exact meet, then $\mathfrak{c}(x_{i})\cap\bigvee_{j}S_{j}\subseteq\mathfrak{c}(x)$ implies that $\mathfrak{c}(x_{i})\cap S_{j}\subseteq\mathfrak{c}(x)$ for all $j$’s, by Lemma 2.4. Therefore, for all $j$’s, $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S_{j}\subseteq\mathfrak{c}(x)$, and the result follows again by linearity. To see that it contains all closed sublocales, consider that if $\mathfrak{c}(x_{i})\cap\mathfrak{c}(y)\subseteq\mathfrak{c}(x)$ then $\mathfrak{c}(x_{i}\vee y)\subseteq\mathfrak{c}(x)$, that is $x\leq x_{i}\vee y$, and so $x\leq\bigwedge_{i}x_{i}\vee y$, by exactness, and this is equivalent to $\mathfrak{c}(\bigwedge_{i}x_{i})\cap\mathfrak{c}(y)\subseteq\mathfrak{c}(x)$. Finally, for open sublocales, we notice that $\mathfrak{c}(x_{i})\cap\mathfrak{o}(y)\subseteq\mathfrak{c}(x)$ means $\mathfrak{c}(x_{i})\subseteq\mathfrak{c}(y)\vee\mathfrak{c}(x)$, and this, by exactness, means $\mathfrak{c}(\bigwedge_{i}x_{i})\subseteq\mathfrak{c}(y)\vee\mathfrak{c}(x)$, that is $\mathfrak{c}(\bigwedge_{i}x_{i})\cap\mathfrak{o}(y)\subseteq\mathfrak{c}(x)$, as desired. For the third part, consider a covered prime $p\in L$ and suppose that, for an exact meet $\bigwedge_{i}x_{i}\in L$, $\mathfrak{c}(x_{i})\cap\mathfrak{b}(p)\subseteq\mathfrak{c}(x)$. This means that $\mathfrak{b}(p)\subseteq\mathfrak{c}(x)\vee\mathfrak{o}(x_{i})$ for all $i$’s. Using the properties of prime elements in Lemma 2.5, we obtain that either $x\leq p$ or $x_{i}\nleq p$ for all $i\in I$. In the first case, $\mathfrak{b}(p)\subseteq\mathfrak{c}(x)$, and the desired result follows. In the second case, $\bigwedge_{i}(x_{i}\vee p)=\bigwedge_{i}x_{i}\vee p\neq p$, by exactness and coveredness, and so $\bigwedge_{i}x_{i}\nleq p$, from which the desired claim follows. ∎
$D$-sublocales, introduced in [arrieta21], are those sublocales such that the corresponding surjection is in $\mathbf{Frm}_{D}$. These may be seen as a (D-)spatial versions of exact sublocales.
Proposition 7.17.
A D-spatial sublocale is exact if and only if it is a D-sublocale.
Proof.
D-spatial sublocales coincide with joins of two-elements sublocales from covered primes. Consider a D-spatial sublocale $\bigvee_{i}\mathfrak{b}(p_{i})$, for $p_{i}\in\mathsf{pt}_{D}(L)$. If this is exact it is also a D-sublocale, by Lemma 7.10. If it is a D-sublocale, then it is exact by Lemma 7.11. ∎
Corollary 7.18.
A two-element sublocale $\mathfrak{b}(p)$ is exact if and only if $p$ is covered.
Lemma 7.19.
If a subcollection $\mathcal{S}\subseteq\mathsf{S}(L)$ is closed under joins and is stable under the operation $-\cap\mathfrak{c}(x)$ and $-\cap\mathfrak{o}(x)$ for all $x\in L$, then it is a subcolocale.
Proof.
Suppose that $\mathcal{S}\subseteq\mathsf{S}(L)$ is closed under all joins and stable under the two operations above. For it to be a subcolocale, it suffices to show that if $S\in\mathcal{S}$ and $T\in\mathsf{S}(L)$ then $S{\setminus}T\in\mathcal{S}$. Every sublocale of $L$ is of the form $\bigcap_{i}\mathfrak{o}(x_{i})\vee\mathfrak{c}(y_{i})$, and $S{\setminus}\bigcap_{i}\mathfrak{o}(x_{i})\vee\mathfrak{c}(y_{i})=\bigvee_{i}(S{\setminus}(\mathfrak{o}(x_{i})\vee\mathfrak{c}(y_{i})))$. Then, for $\mathcal{S}$ to be a subcolocale it suffices for it to be stable under $-{\setminus}(\mathfrak{o}(x)\vee\mathfrak{c}(y))$. If $\mathcal{S}$ is as required, and $S\in\mathcal{S}$, and $x,y\in L$, $S\cap\mathfrak{c}(x)\cap\mathfrak{o}(y)=S{\setminus}(\mathfrak{o}(x)\vee\mathfrak{c}(y))\in\mathcal{S}$. ∎
Theorem 7.20.
The inclusion $\mathsf{S}_{\mathcal{E}}(L)\subseteq\mathsf{S}(L)$ is a subcolocale inclusion.
Proof.
By Lemma 7.19, it suffices to show that the collection is closed under all joins and stable under $-\cap\mathfrak{c}(x)$ and $-\cap\mathfrak{o}(x)$ for all $x\in L$. The first claim follows from Proposition 7.15. For the second, suppose that $y\in L$. Suppose that $S$ is exact. We show that $S\cap\mathfrak{o}(y)$ is exact. If, for exact $\bigwedge_{i}x_{i}$, $\mathfrak{c}(x_{i})\cap S\cap\mathfrak{o}(y)\subseteq\mathfrak{c}(x)$, then $\mathfrak{c}(x_{i})\cap S\subseteq\mathfrak{c}(x)\vee\mathfrak{c}(y)=\mathfrak{c}(x\wedge y)$, and so by hypothesis $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S\subseteq\mathfrak{c}(x\wedge y)$, that is $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S\cap\mathfrak{o}(y)\subseteq\mathfrak{c}(x)$. Let us show that $S\cap\mathfrak{c}(y)$ is exact. For exact $\bigwedge_{i}x_{i}$, if $\mathfrak{c}(x_{i})\cap S\cap\mathfrak{c}(y)\subseteq\mathfrak{c}(x)$ then $\mathfrak{c}(x_{i}\vee y)\cap S\subseteq\mathfrak{c}(x)$, and since $\bigwedge_{i}(x_{i}\vee y)$ is exact by Lemma 7.13, and by exactness of $\bigwedge_{i}x_{i}$, this implies that $\mathfrak{c}(\bigwedge_{i}x_{i}\vee y)\cap S\subseteq\mathfrak{c}(x)$, that is $\mathfrak{c}(\bigwedge_{i}x_{i})\cap S\cap\mathfrak{c}(y)\subseteq\mathfrak{c}(x)$. ∎
For every frame $L$, there are subcolocale inclusions $\mathsf{S}_{\mathcal{E}}(L)\subseteq\mathsf{S}_{D}(L)\subseteq\mathsf{S}(L)$. We do not know, yet, how to characterize frames for which $\mathsf{S}_{\mathcal{E}}(L)=\mathsf{S}_{D}(L)$, or those such that $\mathsf{S}_{\mathcal{E}}(L)=\mathsf{S}(L)$, and leave this as an open question.
[…]